OpenStudy (anonymous):

solve the euation for x: x ^{3}-9x ^{2}-4x+36=0

jhonyy9 (jhonyy9):

x*3 - 9x*2 -4x +36 =0 x(x*2 -4) -9(x*2 -4) =0 (x-9)(x*2 -4)=0 x-9=0 x=9 x*2 -4=0 x*2 =4 x_1,_2 =+/- 2